### v1

# An alternating low-rank projection approach for partial differential equations with random inputs

Abstract
It is known that standard stochastic Galerkin methods face challenges when solving partial differential equations (PDEs) with random inputs. These challenges are typically attributed to the large number of required physical basis functions and stochastic basis functions. Therefore, it becomes crucial to select effective basis functions to properly reduce the dimensionality of both the physical and stochastic approximation spaces. In this study, our focus is on the stochastic Galerkin approximation associated with generalized polynomial chaos (gPC). We delve into the low-rank approximation of the quasimatrix, whose columns represent the coefficients in the gPC expansions of the solution. We conduct an investigation into the singular value decomposition (SVD) of this quasimatrix, proposing a strategy to identify the rank required for a desired accuracy. Subsequently, we introduce both a simultaneous low-rank projection approach and an alternating low-rank projection approach to compute the low-rank approximation of the solution for PDEs with random inputs. Numerical results demonstrate the efficiency of our proposed methods for both diffusion and Helmholtz problems.
keywords:
PDEs with random inputs, low-rank approximation, quasimatrix, generalized polynomial chaos, stochastic Galerkin method.
MSC:
35B30, 35R60, 65C30, 65D40.
[…]
1 Introduction
Over the past few decades, there has been rapid development in efficient numerical methods for solving partial differential equations (PDEs) with random inputs. This surge in interest has been driven by the necessity of performing uncertainty quantification when modeling practical problems, such as groundwater flow and acoustic scattering. The sources of uncertainty in these problems typically arise from lack of knowledge or accurate measurements of realistic model parameters, such as permeability coefficients and refraction coefficients [1, 2, 3, 4, 5].
Among uncertainty quantification approaches, this work specifically focuses on the stochastic Galerkin methods, which have been demonstrated to be effective in various disciplines [6, 7, 8, 9, 10]. In the context of stochastic Galerkin methods, this work focus on the generalized polynomial chaos (gPC) methods to discretize the stochastic parameter space [1, 11], and employ finite element methods to discretize the physical space [12].
Designing efficient solvers is a crucial and challenging problem when dealing with the stochastic Galerkin methods, as it often leads to large coupled linear systems. Since the stochastic Galerkin methods typically result in linear systems formulated in the Kronecker form [13, 14, 15], various iterative algorithms that exploit the Kronecker structure of linear systems to reduce computational costs have been extensively studied in recent years [16, 17, 18, 19, 20]. Additionally, it has been shown that the desired stochastic Galerkin solution can be approximated through low-rank approximation, which further reduces computational costs [21].
In this work, we delve into low-rank approximations for solving PDEs with random inputs, and propose an alternating low-rank projection (AltLRP) approach. To achieve this, we first introduce the concept and properties of the quasimatrix [22, 23, 24]. Then, we investigate the singular value decomposition (SVD) of quasimatrices associated with the truncated gPC expansions of the solutions. The relationship between the singular values of these quasimatrices is explored in detail, accompanied by a thorough theoretical analysis. Building upon these findings, a systematic strategy is proposed to identify the rank for a desired accuracy, and a simultaneous low-rank projection (SimLRP) approach is developed, where reduced basis functions of the stochastic and physical spaces are constructed simultaneously. Based on SimLRP, our AltLRP approach is proposed to refine the solution, with the outcome of SimLRP serving as the initial guess for AltLRP. The effectiveness of AltLRP is illustrated through numerical studies.
The outline of this work is as follows. In Section 2, we describe the PDE formulation considered in this study and introduce the stochastic Galerkin method. In Section 3, our main theoretical analysis concerning the relationship between the singular values of quasimatrices associated with gPC is discussed, and our AltLRP approach is presented. Numerical results are discussed in Section 4, and Section 5 concludes the paper.
2 Problem setting and stochastic Galerkin method
This section describes the problems considered in this study, revisits the stochastic Galerkin method, and presents the resulting linear system from the stochastic Galerkin method.
2.1 Problem setting
Let $D$ be a bounded and connected physical domain in $\mathbf{R}^{d}~{}(d=2,3)$ with a polygonal boundary denoted as $\partial D$ and $\bm{x}\in\mathbf{R}^{d}$ represent a physical variable. Let $\bm{\xi}=(\xi_{1},\ldots,\xi_{N})$ be a random vector of dimension $N$, where the image of $\xi_{i}$ is denoted by $\Gamma_{i}$, and the probability density function of $\xi_{i}$ is denoted by $\rho_{i}(\xi_{i})$. We further assume that the components of $\bm{\xi}$, i.e., $\xi_{1},\ldots,\xi_{N}$ are mutually independent, the image of $\bm{\xi}$ is then given by $\Gamma=\Gamma_{1}\times\cdots\times\Gamma_{N}$, and the probability density function of $\bm{\xi}$ is given by $\rho(\bm{\xi})=\prod_{i=1}^{N}\rho_{i}(\xi_{i})$. In this work, we consider the following partial differential equations (PDEs) with random inputs, i.e.,
$\begin{cases}\mathcal{L}(\bm{x},\bm{\xi},u(\bm{x},\bm{\xi}))=f(\bm{x})\ &\forall~{}\bm{x}\in D,\\ \mathcal{B}(\bm{x},\bm{\xi},u(\bm{x},\bm{\xi}))=g(\bm{x})\ &\forall~{}\bm{x}\in\partial D,\end{cases}$ (1)
where $\mathcal{L}$ is a linear partial differential operator with respect to the physical variable, and $\mathcal{B}$ is a boundary operator. Both operators may have random coefficients. The source function is denoted by $f(\bm{x})$, and $g(\bm{x})$ specifies the boundary conditions. Additionally, we assume that $\mathcal{L}$ and $\mathcal{B}$ are affinely dependent on the random inputs, i.e.,
$\mathcal{L}(\bm{x},\bm{\xi},u(\bm{x},\bm{\xi}))=\sum_{i=1}^{K}\mathrm{\Theta}_{\mathcal{L}}^{(i)}(\bm{\xi})\mathcal{L}_{i}(\bm{x},u(\bm{x},\bm{\xi})),$ (2)
$\mathcal{B}(\bm{x},\bm{\xi},u(\bm{x},\bm{\xi}))=\sum_{i=1}^{K}\mathrm{\Theta}_{\mathcal{B}}^{(i)}(\bm{\xi})\mathcal{B}_{i}(\bm{x},u(\bm{x},\bm{\xi})),$ (3)
where $\{\mathcal{L}_{i}\}_{i=1}^{K}$ are parameter-independent linear differential operators, and $\{\mathcal{B}_{i}\}_{i=1}^{K}$ are parameter-independent boundary operators. Both $\mathrm{\Theta}_{\mathcal{L}}^{(i)}(\bm{\xi})$ and $\mathrm{\Theta}_{\mathcal{B}}^{(i)}(\bm{\xi})$ take values in $\mathbf{R}$ for $i=1,\ldots,K$.
It is of interest to design a surrogate model for the problem (1) or calculate statistics of the stochastic solution $u(\bm{x},\bm{\xi})$, such as the mean and the variance. For simplicity of the presentation, we consider problems that satisfy homogeneous Dirichlet boundary conditions. However, it is noteworthy that the approach we present can be readily extended to other arbitrary well-posed boundary conditions.
2.2 Stochastic Galerkin method
We provide a concise overview of the stochastic Galerkin method, as detailed in [6, 25]. To facilitate our discussion, we begin by introducing the variational formulation of (1). In order to do this, it is necessary to establish some key notations. Firstly, we define the Hilbert spaces $L^{2}(D)$ and $L^{2}_{\rho}(\Gamma)$ as follows:
$L^{2}(D)$ ::=:= $\left\{\psi(\bm{x}):D\to\mathbf{R}\ \bigg{|}\ \int_{D}\psi^{*}(\bm{x})\psi(\bm{x})\,\mathrm{d}\bm{x}<\infty\right\},$
$L_{\rho}^{2}(\Gamma)$ ::=:= $\left\{\phi(\bm{\xi}):\Gamma\to\mathbf{R}\ \bigg{|}\ \int_{\Gamma}\rho(\bm{\xi})\phi^{*}(\bm{\xi})\phi(\bm{\xi})\,\mathrm{d}\bm{\xi}<\infty\right\},$
which are equipped with the inner products
$\langle\psi(\bm{x}),\widetilde{\psi}(\bm{x})\rangle_{L^{2}}:=\int_{D}\widetilde{\psi}^{*}(\bm{x})\psi(\bm{x})\,\mathrm{d}\bm{x},\ \langle\phi(\bm{\xi}),\widetilde{\phi}(\bm{\xi})\rangle_{L^{2}_{\rho}}:=\int_{\Gamma}\rho(\bm{\xi})\widetilde{\phi}^{*}(\bm{\xi})\phi(\bm{\xi})\,\mathrm{d}\bm{\xi}.$
Following [25], we define the tensor space of $L^{2}(D)$ and $L^{2}_{\rho}(\Gamma)$ as
$L^{2}(D)\otimes L^{2}_{\rho}(\Gamma):=\left\{\omega(\bm{x},\bm{\xi})\bigg{|}\omega(\bm{x},\bm{\xi})=\sum_{i=1}^{n}\psi_{i}(\bm{x})\phi_{i}(\bm{\xi}),\psi_{i}(\bm{x})\in L^{2}(D),\,\phi_{i}(\bm{\xi})\in L^{2}_{\rho}(\Gamma),n\in\mathbf{N}^{+}\right\},$
which is equipped with the inner product
$\left\langle\omega(\bm{x},\bm{\xi}),\widetilde{\omega}(\bm{x},\bm{\xi})\right\rangle_{L^{2}\otimes L_{\rho}^{2}}=\int_{\Gamma}\int_{D}\rho(\bm{\xi})\widetilde{\omega}^{*}(\bm{x},\bm{\xi})\omega(\bm{x},\bm{\xi})\mathrm{d}\bm{x}\mathrm{d}\bm{\xi}.$
In general, the variational form of (1) can be expressed as follows: find $u$ in $W=X(D)\otimes L^{2}_{\rho}(\Gamma)$ such that
$\mathcal{A}(u,\omega)=\mathcal{F}(\omega),\ \forall~{}\omega\in W,$ (4)
where $X(D)$ is an appropriate subspace of $L^{2}(D)$ associated with $D$, and
$\mathcal{A}(u,\omega):=\left\langle\mathcal{L}(\bm{x},\bm{\xi},u(\bm{x},\bm{\xi})),\omega(\bm{x},\bm{\xi})\right\rangle_{L^{2}\otimes L^{2}_{\rho}},\ \mathcal{F}(\omega):=\left\langle f(\bm{x}),\omega(\bm{x},\bm{\xi})\right\rangle_{L^{2}\otimes L^{2}_{\rho}}.$
The stochastic Galerkin method aims to approximate (4) within a finite-dimensional subspace of $W$. This subspace is defined as
$W_{h}^{p}:=X_{h}\otimes S_{p}:={\rm span}\left\{\psi_{s}(\bm{x})\phi_{j}(\bm{\xi})\left|\psi_{s}\in X_{h},\phi_{j}\in S_{p}\right.\right\},$
where
$\begin{split}X_{h}={\rm span}\left\{\psi_{s}(\bm{x})\right\}_{s=1}^{n_{x}}\subseteq X(D),\ S_{p}={\rm span}\left\{\phi_{j}(\bm{\xi})\right\}_{j=1}^{n_{\xi}}\subseteq L^{2}_{\rho}(\Gamma).\end{split}$
In this context, $\{\psi_{s}(\bm{x})\}_{s=1}^{n_{x}}$ represents physical basis functions, while $\{\phi_{j}(\bm{\xi})\}_{j=1}^{n_{\xi}}$ represents stochastic basis functions.
In this study, we employ quadrilateral elements and piecewise bilinear basis functions to discretize the physical domain, while for the stochastic domain, we utilize the generalized polynomial chaos basis functions [1]. Note that in this situation, $\{\phi_{j}(\bm{\xi})\}_{j=1}^{n_{\xi}}$ is orthonormal in $L^{2}_{\rho}(\Gamma)$.
Assuming that $u^{{\scriptsize\mbox{ap}}}(\bm{x},\bm{\xi})$ is defined as
$u^{{\scriptsize\mbox{ap}}}(\bm{x},\bm{\xi}):=\sum_{j=1}^{n_{\xi}}\sum_{s=1}^{n_{x}}u_{sj}\psi_{s}(\bm{x})\phi_{j}(\bm{\xi}).$ (5)
Since $\mathcal{L}$ is affinely dependent on the random inputs (see (2)), we substitute (5) into (4) to obtain [15],
$\left(\sum_{i=1}^{K}\bm{G}_{i}\otimes\bm{A}_{i}\right)\bm{u}=\bm{h}\otimes\bm{f},$ (6)
where $\{\bm{G}_{i}\}_{i=1}^{K}$ are matrices of size $n_{\xi}\times n_{\xi}$, and $\bm{h}$ is a column vector of length $n_{\xi}$. They are defined via
$\bm{G}_{i}(j,k)=\langle\mathrm{\Theta}_{\mathcal{L}}^{(i)}(\bm{\xi})\phi_{j}(\bm{\xi}),\phi_{k}(\bm{\xi})\rangle_{L^{2}_{\rho}},\ \bm{h}(k)=\langle 1,\phi_{k}(\bm{\xi})\rangle_{L^{2}_{\rho}}.$ (7)
The matrices $\bm{A}_{i}$ and the vector $\bm{f}$ in (6) are defined through
$\bm{A}_{i}(s,t)=\left\langle\mathcal{L}_{i}\psi_{s},\psi_{t}\right\rangle_{L^{2}},\ \bm{f}(s)=\left\langle f,\psi_{s}\right\rangle_{L^{2}},$ (8)
where $s=1,\ldots n_{x}$ and $t=1,\ldots,n_{x}$. The vector $\bm{u}$ in (6) is a column vector of length $n_{x}\times n_{\xi}$, and is defined by
$\bm{u}=\begin{bmatrix}\bm{u}_{1}\\ \vdots\\ \bm{u}_{n_{\xi}}\end{bmatrix},\ \mbox{where}\ \bm{u}_{j}=\begin{bmatrix}{u}_{1j}\\ \vdots\\ {u}_{n_{x}j}\end{bmatrix},\ j=1,\ldots,n_{\xi}.$
The linear system (6) can be rewritten in a matrix form
$\sum_{i=1}^{K}\bm{A}_{i}\bm{U}\bm{G}_{i}^{\mathrm{T}}=\bm{f}\bm{h}^{\mathrm{T}},$ (9)
where $\bm{U}=[\bm{u}_{1},\ldots,\bm{u}_{n_{\xi}}]$.
3 Alternating low-rank projection approach
In this section, we introduce an alternating low-rank projection (AltLRP) approach for PDEs with random inputs. To this end, we first present the low-rank approximation of a quasimatrix and investigate the singular value decomposition (SVD) of the quasimatrix. The columns of this quasimatrix represent the coefficients of the truncated generalized polynomial chaos expansions corresponding to the solution for (1). Additionally, we explore the SVD of another quasimatrix, where the columns represent the values of the truncated generalized polynomial chaos expansions at random samples of random inputs (i.e., $\bm{\xi}$). The relationship between the singular values of these two quasimatrices is also examined.
Based on these findings, we propose a simultaneous low-rank projection (SimLRP) approach for PDEs with random inputs, in which the reduced basis functions of the stochastic and physical spaces are constructed simultaneously. To further improve the performance, we introduce an alternating low-rank projection (AltLRP) approach, in which low-rank projections in the stochastic and physical spaces are applied alternately. AltLRP capitalizes on the outcomes of SimLRP by employing them as an initial guess, leveraging pre-established insights to facilitate a refined and robust convergence toward accurate solutions.
3.1 The low-rank approximation of a quasimatrix
Let us consider the generalized polynomial chaos expansion of $u(\bm{x},\bm{\xi})$, i.e., the solution to (1). Suppose that the generalized polynomial chaos expansion of $u(\bm{x},\bm{\xi})$ is given by
$u(\bm{x},\bm{\xi})=\sum_{j=1}^{\infty}u_{j}(\bm{x})\phi_{j}(\bm{\xi}),$
and we truncate this expansion at a gPC order $p$, i.e.,
$u^{p}(\bm{x},\bm{\xi})=\sum_{j=1}^{n_{\xi}}u_{j}(\bm{x})\phi_{j}(\bm{\xi}),\ \ (\bm{x},\bm{\xi})\in D\times\Gamma,$ (10)
where $n_{\xi}=(N+p)!/(N!p!)$. Rewrite (10) in an analogues matrix form
$u^{p}(\bm{x},\bm{\xi})=\mathbb{U}^{p}_{D\times n_{\xi}}(\mathbb{S}^{p}_{\Gamma\times n_{\xi}})^{*},$ (11)
where
$\mathbb{U}^{p}_{D\times n_{\xi}}=\left[\begin{array}[]{llll}\Big{|}&\Big{|}&&\Big{|}\\ u_{1}&u_{2}&\ldots&u_{n_{\xi}}\\ \Big{|}&\Big{|}&&\Big{|}\end{array}\right],\ \mathbb{S}^{p}_{\Gamma\times n_{\xi}}=\left[\begin{array}[]{llll}\Big{|}&\Big{|}&&\Big{|}\\ \phi_{1}^{*}&\phi_{2}^{*}&\ldots&\phi_{n_{\xi}}^{*}\\ \Big{|}&\Big{|}&&\Big{|}\end{array}\right].$
We refer to $\mathbb{U}^{p}_{D\times n_{\xi}}$ and $\mathbb{S}^{p}_{\Gamma\times n_{\xi}}$ as $D\times n_{\xi}$ and $\Gamma\times n_{\xi}$ quasimatrices, respectively, since one index of the rectangular matrix becomes continuous while the other remains discrete [24]. We use the superscript $\mathrm{T}$ to denote the transpose and * to denote the conjugate transpose of a quasimatrix. More precisely, suppose that
$\mathbb{A}=\left[\begin{array}[]{llll}\Big{|}&\Big{|}&&\Big{|}\\ a_{1}&a_{2}&\ldots&a_{n}\\ \Big{|}&\Big{|}&&\Big{|}\end{array}\right]$
is a $D\times n$ quasimatrix, then
$\mathbb{A}^{\mathrm{T}}=\left[\begin{array}[]{ccc}\text{---}&a_{1}&\text{---}\\ \text{---}&a_{2}&\text{---}\\ &\vdots&\\ \text{---}&a_{n}&\text{---}\end{array}\right],\ \mathbb{A}^{*}=\left[\begin{array}[]{ccc}\text{---}&a_{1}^{*}&\text{---}\\ \text{---}&a_{2}^{*}&\text{---}\\ &\vdots&\\ \text{---}&a_{n}^{*}&\text{---}\end{array}\right].$
Thus $(\mathbb{S}^{p}_{\Gamma\times n_{\xi}})^{*}$ in (11) is given by
$(\mathbb{S}^{p}_{\Gamma\times n_{\xi}})^{*}=\left[\begin{array}[]{ccc}\text{---}&\phi_{1}&\text{---}\\ \text{---}&\phi_{2}&\text{---}\\ &\vdots&\\ \text{---}&\phi_{n_{\xi}}&\text{---}\end{array}\right].$
Note that if $a_{i}(i=1,\ldots,n)$ are real functions, then $\mathbb{A}^{\mathrm{T}}=\mathbb{A}^{*}$. The product of $\mathbb{A}^{*}$ and $\mathbb{A}$ is defined as
$\mathbb{A}^{*}\mathbb{A}=\left[\begin{array}[]{cccc}a_{1}^{*}a_{1}&a_{1}^{*}a_{2}&\ldots&a_{1}^{*}a_{n}\\ a_{2}^{*}a_{1}&a_{2}^{*}a_{2}&\ldots&a_{2}^{*}a_{n}\\ \vdots&\vdots&\ldots&\vdots\\ a_{n}^{*}a_{1}&a_{n}^{*}a_{2}&\ldots&a_{n}^{*}a_{n}\\ \end{array}\right].$
For a matrix, QR factorization exists. The QR factorization of a $D\times n$ quasimatrix $\mathbb{A}$ is a straightforward extension of the matrix case. Following [24], we first give the definition of the QR factorization of a quasimatrix.
Definition 3.1 (QR factorization of a quasimatrix).
Let $\mathbb{A}$ be a $D\times n$ quasimatrix. A QR factorization of $\mathbb{A}$ is a factorization $\mathbb{A}=\mathbb{Q}\bm{R}$, where $\mathbb{Q}$ is a $D\times n$ quasimatrix with orthonormal columns and $\bm{R}$ is an $n\times n$ upper triangular matrix with non-negative real numbers on the diagonal.
The QR factorization of a $D\times n$ quasimatrix has the following fundamental properties [22, 23, 24].
Theorem 3.1.
Every $D\times n$ quasimatrix has a QR factorization, which can be calculated by Gram–Schmidt orthogonalization. If the columns of $\mathbb{A}$ are linearly independent, the QR factorization is unique.
Proof.
The existence is evident through the Gram–Schmidt process, and we focus here on providing the proof of uniqueness [^1] The proof is available in [22]; however, we are unable to access the literature in electronic or hard copy form.. Suppose that
$\mathbb{A}=\mathbb{Q}_{1}\bm{R}_{1}=\mathbb{Q}_{2}\bm{R}_{2}.$
Since the columns of $\mathbb{A}$ are linearly independent, $\bm{R}_{1}$ and $\bm{R}_{2}$ are invertible, and the diagonal elements of $\bm{R}_{1}$ and $\bm{R}_{2}$ are positive. Thus we have
$\mathbb{Q}_{1}=\mathbb{Q}_{2}\bm{R}_{2}\bm{R}_{1}^{-1},$
which implies
$\mathbb{Q}_{1}^{*}\mathbb{Q}_{1}=(\bm{R}_{2}\bm{R}_{1}^{-1})^{*}\mathbb{Q}_{2}^{*}\mathbb{Q}_{2}\bm{R}_{2}\bm{R}_{1}^{-1}.$
Since the columns of $\mathbb{Q}_{1}$ and $\mathbb{Q}_{2}$ are orthonormal, by integration over $D$, we obtain
$\bm{E}=(\bm{R}_{2}\bm{R}_{1}^{-1})^{*}\bm{R}_{2}\bm{R}_{1}^{-1}$ (12)
Note that $\bm{R}_{1}$ and $\bm{R}_{2}$ are upper triangular matrix, and thus $\bm{R}^{-1}$ and $\bm{R}_{2}\bm{R}_{1}^{-1}$ are also upper triangular matrices. The diagonal elements of $\bm{R}_{2}\bm{R}_{1}^{-1}$ are given by $\bm{R}_{2}(i,i)/\bm{R}_{1}(i,i)$.
Suppose that
$\bm{R}_{2}\bm{R}_{1}^{-1}=\begin{bmatrix}r_{11}&r_{12}&\ldots&r_{1n}\\ 0&r_{22}&\ldots&r_{2n}\\ 0&0&\ddots&\vdots\\ 0&0&\ldots&r_{nn}\end{bmatrix},\ (\bm{R}_{2}\bm{R}_{1}^{-1})^{*}=\begin{bmatrix}r_{11}^{*}&0&\ldots&0\\ r_{12}^{*}&r_{22}^{*}&\ldots&0\\ \vdots&\vdots&\ddots&\vdots\\ r_{1n}^{*}&r_{2n}^{*}&\ldots&r_{nn}^{*}\end{bmatrix}.$
Since $r_{ii}=\bm{R}_{2}(i,i)/\bm{R}_{1}(i,i)>0$, by (12), we have $\bm{R}_{2}\bm{R}_{1}^{-1}=\bm{E}$, and thus
$\bm{R}_{1}=\bm{R}_{2},\ \mathbb{Q}_{1}=\mathbb{Q}_{2}.$
Definition 3.2 (SVD decomposition of a quasimatrix).
Let $\mathbb{A}$ be a $D\times n$ quasimatrix. An SVD of $\mathbb{A}$ is a factorization $\mathbb{A}=\mathbb{U}\bm{\Sigma V}^{*}$, where $\mathbb{U}$ is a $D\times n$ quasimatrix with orthonormal columns, $\bm{\Sigma}$ is an $n\times n$ diagonal matrix with diagonal entries $\sigma_{1}\geq\sigma_{2}\geq\ldots\sigma_{n}\geq 0$ and $\bm{V}$ is an $n\times n$ unitary matrix.
Theorem 3.2.
Every $D\times n$ quasimatrix has an SVD, which can be calculated by computing a QR decomposition $\mathbb{A}=\mathbb{Q}\bm{R}$ followed by a matrix SVD of the triangular factor, $\bm{R}=\bm{U}_{1}\bm{\Sigma V}^{*}$; an SVD of $\mathbb{A}$ is then obtained as $(\mathbb{Q}\bm{U}_{1})\bm{\Sigma V}^{*}$. The singular values are unique, and the singular vectors corresponding to simple singular values are also unique up to complex signs. The rank $r$ of $\mathbb{A}$ is the number of nonzero singular values.
Proof.
The existence is obvious, and the proof of other conclusions can be found in [23]. ∎
Theorem 3.3.
Suppose that $\bm{\xi}_{1},\ldots\bm{\xi}_{n}$ are random samplings from $\bm{\xi}$, and the quasimatrix $\mathbb{U}^{{\mathrm{samp}}}_{D\times n}$ is defined as
$\mathbb{U}^{{\mathrm{samp}}}_{D\times n}=\left[\begin{array}[]{llll}\Big{|}&\Big{|}&&\Big{|}\\ u^{p}(\bm{x},\bm{\xi}_{1})&u^{p}(\bm{x},\bm{\xi}_{2})&\ldots&u^{p}(\bm{x},\bm{\xi}_{n})\\ \Big{|}&\Big{|}&&\Big{|}\end{array}\right].$
Then the number of non-zero singular values of $\mathbb{U}^{{\mathrm{samp}}}_{D\times n}$ is less than or equal to $n_{\xi}$. Denote the first $n_{\xi}$ singular values of $\mathbb{U}^{{\mathrm{samp}}}_{D\times n}$ by $\lambda_{1}\geq\lambda_{2}\geq\ldots\geq\lambda_{n_{\xi}}\geq 0$, then for large $n$ we have
$\sqrt{\sum_{j=1}^{n_{\xi}}\left|\sigma_{j}^{2}-\frac{\lambda_{j}^{2}}{n}\right|^{2}}\left/\sum_{j=1}^{n_{\xi}}\sigma_{j}^{2}\right.\lesssim c\frac{n_{\xi}}{\sqrt{n}}\gamma.$
where $\sigma_{1}\geq\sigma_{2}\geq\ldots\geq\sigma_{n_{\xi}}\geq 0$ are the singular values of $\mathbb{U}^{p}_{D\times n_{\xi}}$, $c$ is a constant and $\gamma$ is a random variable.
Proof.
Suppose that the SVD decomposition of $\mathbb{U}^{p}_{D\times n_{\xi}}$ is given by
$\mathbb{U}^{p}_{D\times n_{\xi}}=\mathbb{U}\bm{\Sigma}\bm{V}^{*},$
then we have
$\mathbb{U}^{{\mathrm{samp}}}_{D\times n}=\mathbb{U}\bm{\Sigma}\bm{V}^{*}\bm{\Phi}^{*},$
where
$\mathbb{U}^{{\mathrm{samp}}}_{D\times n}=\left[\begin{array}[]{llll}\Big{|}&\Big{|}&&\Big{|}\\ u^{p}(\bm{x},\bm{\xi}_{1})&u^{p}(\bm{x},\bm{\xi}_{2})&\ldots&u^{p}(\bm{x},\bm{\xi}_{n})\\ \Big{|}&\Big{|}&&\Big{|}\end{array}\right],\bm{\Phi}^{*}=\left[\begin{array}[]{cccc}\phi_{1}(\bm{\xi}_{1})&\phi_{1}(\bm{\xi}_{2})&\ldots&\phi_{1}(\bm{\xi}_{n})\\ \phi_{2}(\bm{\xi}_{1})&\phi_{2}(\bm{\xi}_{2})&\ldots&\phi_{2}(\bm{\xi}_{n})\\ \vdots&\vdots&\ldots&\vdots\\ \phi_{n_{\xi}}(\bm{\xi}_{1})&\phi_{n_{\xi}}(\bm{\xi}_{n})&\ldots&\phi_{n_{\xi}}(\bm{\xi}_{n})\\ \end{array}\right].$
Since the rank of $\mathbb{U}^{{\mathrm{samp}}}_{D\times n}$ is less than or equal to $n_{\xi}$, the number of non-zero singular values is less than or equal to $n_{\xi}$. The non-zero singular values of $\mathbb{U}^{{\mathrm{samp}}}_{D\times n}$ can be given by the square root of the eigenvalues of  [26]
$\bm{T}:=\bm{\Sigma}\bm{V}^{*}\bm{\Phi}^{*}\bm{\Phi}\bm{V}\bm{\Sigma}^{*}.$ (13)
Now let us consider the eigenvalues of $\bm{T}$. Note that $\{\phi_{1},\ldots,\phi_{n_{\xi}}\}$ is orthonormal, that is
$\left\langle\phi_{i},\phi_{j}\right\rangle_{L^{2}_{\rho}}=\int_{\Gamma}\phi_{j}^{*}(\bm{\xi})\phi_{i}(\bm{\xi})\mathrm{d}\bm{\xi}=\delta_{ij},$
where $\delta_{ij}$ is the Kronecker’s delta function. Thus for large $n$, we have [27]
$\frac{1}{n}\sum_{k=1}^{n}\phi_{i}(\bm{\xi}_{k})\phi^{*}_{j}(\bm{\xi}_{k})-\delta_{ij}\approx c_{ij}\frac{1}{\sqrt{n}}\gamma_{ij},\ i,j=1\ldots,n_{\xi},$
where $c_{ij}$ is a constant, and $\gamma_{ij}$ is a standard normal random variable. This implies that
$\left\|\frac{1}{n}\bm{\Phi}^{*}\bm{\Phi}-\bm{E}\right\|_{F}\lesssim c\frac{n_{\xi}}{\sqrt{n}}\gamma,$
where $c=\max\{c_{ij}\}$, is a constant, and $\gamma=\max\{|\gamma_{ij}|\}$ is a random variable. Suppose that
$\frac{1}{n}\bm{\Phi}^{*}\bm{\Phi}=\bm{E}+\bm{\Delta},\ \mbox{where}\ \|\bm{\Delta}\|_{F}\lesssim c\frac{n_{\xi}}{\sqrt{n}}\gamma.$
Note that
$\begin{split}\bm{T}=&(\sqrt{n}\bm{\Sigma})\bm{V}^{*}\left(\frac{1}{n}\bm{\Phi}^{*}\bm{\Phi}\right)\bm{V}(\sqrt{n}\bm{\Sigma})=(\sqrt{n}\bm{\Sigma})\bm{V}^{*}(\bm{E}+\bm{\Delta})\bm{V}(\sqrt{n}\bm{\Sigma})\\ =&n\bm{\Sigma}^{2}+n(\bm{V}\bm{\Sigma})^{*}\bm{\Delta}(\bm{V}\bm{\Sigma}).\end{split}$
It is clear that the eigenvalues of $n\bm{\Sigma}^{2}$ are $n\sigma_{1}^{2},\ldots,n\sigma_{n_{\xi}}^{2}$, and the eigenvalues of $\bm{T}$ are $\lambda_{1}^{2},\ldots,\lambda_{n_{\xi}}^{2}$. Since $\bm{T}$ is a normal matrix, we have [28, 29, 30]
$\sqrt{\sum_{j=1}^{n_{\xi}}|n\sigma_{j}^{2}-\lambda_{j}^{2}|^{2}}\leq\|n(\bm{V}\bm{\Sigma})^{*}\bm{\Delta}(\bm{V}\bm{\Sigma})\|_{F}.$
Note that
$\begin{split}\|n(\bm{V}\bm{\Sigma})^{*}\bm{\Delta}(\bm{V}\bm{\Sigma})\|_{F}\leq n\|\Delta\|_{F}\|\bm{V}\bm{\Sigma}\|_{F}^{2}=n\|\bm{\Sigma}\|_{F}^{2}\|\bm{\Delta}\|_{F}\lesssim c\sqrt{n}n_{\xi}\|\bm{\Sigma}\|_{F}^{2}\gamma,\end{split}$
which implies
$\sqrt{\sum_{j=1}^{n_{\xi}}|n\sigma_{j}^{2}-\lambda_{j}^{2}|^{2}}\lesssim c\sqrt{n}n_{\xi}\|\bm{\Sigma}\|_{F}^{2}\gamma.$
This further implies
$\sqrt{\sum_{j=1}^{n_{\xi}}\left|\sigma_{j}^{2}-\frac{\lambda_{j}^{2}}{n}\right|^{2}}\lesssim c\frac{n_{\xi}}{\sqrt{n}}\|\bm{\Sigma}\|_{F}^{2}\gamma.$
That is
$\sqrt{\sum_{j=1}^{n_{\xi}}\left|\sigma_{j}^{2}-\frac{\lambda_{j}^{2}}{n}\right|^{2}}\left/\sum_{j=1}^{n_{\xi}}\sigma_{j}^{2}\right.\lesssim c\frac{n_{\xi}}{\sqrt{n}}\gamma.$
Corollary 3.3.1.
Suppose that $\bm{\xi}_{1},\ldots\bm{\xi}_{n}$ are random samplings from $\bm{\xi}$, and the quasimatrix $\widehat{\mathbb{U}}^{{\mathrm{samp}}}_{D\times n}$ is defined as
$\widehat{\mathbb{U}}^{{\mathrm{samp}}}_{D\times n}=\mathbb{U}^{{\mathrm{samp}}}_{D\times n}\mathrm{diag}(\frac{1}{\sqrt{n}},\ldots,\frac{1}{\sqrt{n}}).$
Then the number of non-zero singular values of $\widehat{\mathbb{U}}^{{\mathrm{samp}}}_{D\times n}$ is less than or equal to $n_{\xi}$. Denote the first $n_{\xi}$ singular values of $\widehat{\mathbb{U}}^{{\mathrm{samp}}}_{D\times n}$ by $\widehat{\lambda}_{1}\geq\widehat{\lambda}_{2}\geq\ldots\geq\widehat{\lambda}_{n_{\xi}}\geq 0$, then for large $n$ we have
$\sqrt{\sum_{j=1}^{n_{\xi}}\left|\sigma_{j}^{2}-\widehat{\lambda}_{j}^{2}\right|^{2}}\left/\sum_{j=1}^{n_{\xi}}\sigma_{j}^{2}\right.\lesssim c\frac{n_{\xi}}{\sqrt{n}}\gamma,$
where $\sigma_{1}\geq\sigma_{2}\geq\ldots\geq\sigma_{n_{\xi}}\geq 0$ are the singular values of $\mathbb{U}^{p}_{D\times n_{\xi}}$, $c$ is a constant and $\gamma$ is a random variable.
Proof.
Note that $\widehat{\lambda}_{j}={\lambda_{j}}/{\sqrt{n}}$, the conclusion follows immediately from Theorem 3.3. ∎
3.2 A simultaneous low-rank projection approach
Now let us consider the quasimatrix corresponding to $u^{p}(\bm{x},\bm{\xi})$. Suppose that the SVD decomposition of $\mathbb{U}^{p}_{D\times n_{\xi}}$ is given by
$\mathbb{U}^{p}_{D\times n_{\xi}}={\mathbb{U}}{\bm{\Sigma}}{\bm{V}}^{*},$
[…]
$u^{p}(\bm{x},\bm{\xi})={\mathbb{U}}{\bm{\Sigma}}{\bm{V}}^{*}(\mathbb{S}^{p}_{\Gamma\times n_{\xi}})^{*},$ (14)
where $(\mathbb{S}^{p}_{\Gamma\times n_{\xi}})^{*}$ is an $n_{\xi}\times\Gamma$ quasimatrix with orthonormal rows $\phi_{1},\ldots,\phi_{n_{\xi}}$. Since ${\bm{V}}$ is a unitary matrix, the rows of the quasimatrix $\bm{\mathbb{S}}^{*}={\bm{V}}^{*}(\mathbb{S}^{p}_{\Gamma\times n_{\xi}})^{*}$ are also orthonormal. Denote the rows of $\bm{\mathbb{S}}^{*}$ as $\mathring{\phi}_{1},\ldots,\mathring{\phi}_{n_{\xi}}$, we can then rewrite (14) as
[…]
For any $k$, we can consider the partial sum
[…]
It is easy to verify that the error of rank $k$ approximation of $u^{p}(\bm{x},\bm{\xi})$ is given by
$\left\|u^{p}(\bm{x},\bm{\xi})-u^{p}_{k}(\bm{x},\bm{\xi})\right\|_{L^{2}\otimes L^{2}_{\rho}}=\left(\sum_{j=k+1}^{n_{\xi}}{\sigma}_{j}^{2}\right)^{1/2}.$
If we use
[…]
as the basis for the physical and stochastic subspaces respectively, we may result in a much smaller linear system compared to (6). However, since the $D\times n_{\xi}$ quasimatrix $\mathbb{U}^{p}_{D\times n_{\xi}}$ cannot be determined in advance, some adjustments are necessary to adapt this idea to design a solver for (1).
It is worth noting that the accuracy of the rank $k$ approximation of $u^{p}(\bm{x},\bm{\xi})$ depends only on the quasimatrix $\mathbb{U}^{p}_{D\times n_{\xi}}$. In practical implementations, quadrature formulas are typically used to compute the integral in the Gram–Schmidt orthogonalization process for the QR factorization of a quasimatrix. This inspires us to construct the basis functions of the stochastic space and to determine the rank $k$ as follows:
First, we compute a numerical solution on a coarse grid of the physical space. If the coefficients of the gPC basis functions are denoted by $u_{1}^{{\mathrm{c}}}(\bm{x}),\ldots,u_{n_{\xi}}^{{\mathrm{c}}}(\bm{x})$, the solution on the coarse grid of the physical space can be expressed as:
$u^{{\mathrm{c}}}(\bm{x},\bm{\xi})=\mathbb{U}^{{\mathrm{c}}}_{D\times n_{\xi}}(\mathbb{S}^{p}_{\Gamma\times n_{\xi}})^{*},$
where
$\mathbb{U}^{{\mathrm{c}}}_{D\times n_{\xi}}=\left[\begin{array}[]{llll}\Big{|}&\Big{|}&&\Big{|}\\ u_{1}^{{\mathrm{c}}}&u_{2}^{{\mathrm{c}}}&\ldots&u_{n_{\xi}}^{{\mathrm{c}}}\\ \Big{|}&\Big{|}&&\Big{|}\end{array}\right].$
Next, we perform the SVD decomposition of $\mathbb{U}^{{\mathrm{c}}}_{D\times n_{\xi}}$:
$\mathbb{U}^{{\mathrm{c}}}_{D\times n_{\xi}}=\widehat{\mathbb{U}}\widehat{\bm{\Sigma}}\widehat{\bm{V}}^{*},$ (16)
where $\widehat{\mathbb{U}}$ is a $D\times n_{\xi}$ quasimatrix and $\widehat{\bm{\Sigma}}$ is a diagonal matrix. The diagonal elements of $\widehat{\bm{\Sigma}}$, denoted by $\widehat{\sigma}_{1},\ldots,\widehat{\sigma}_{n_{\xi}}$, approximate $\sigma_{1},\ldots,\sigma_{n_{\xi}}$ form (15). Defining
$\widehat{\mathbb{S}}^{*}=\widehat{\bm{V}}^{*}(\mathbb{S}^{p}_{\Gamma\times n_{\xi}})^{*},$ (17)
we note that $\widehat{\mathbb{S}}^{*}$ is a $\Gamma\times n_{\xi}$ quasimatrix with orthonormal rows. Denoting the rows of $\widehat{\mathbb{S}}^{*}$ by $\widehat{\phi}_{1}(\bm{\xi}),\ldots,\widehat{\phi}_{n_{\xi}}(\bm{\xi})$, the low-rank basis functions of stochastic space is given by:
$S^{(k)}=\{\widehat{\phi}_{1},\ldots,\widehat{\phi}_{k}\},$
where $k$ is the smallest number such that
$\left(\sum_{j=k+1}^{n_{\xi}}(\widehat{\sigma}_{j})^{2}\right)^{1/2}\left/\left(\sum_{j=1}^{n_{\xi}}(\widehat{\sigma}_{j})^{2}\right)^{1/2}\right.\leq\mathtt{RelTol}.$ (18)
Here, $\mathtt{RelTol}$ is the desired accuracy.
For the low-rank basis functions of the physical space, choosing the columns of the quasimatrix $\widehat{\mathbb{U}}$ becomes less effective when the mesh grid is refined. Instead, we construct the low-rank basis functions of the physical space based on Theorem 3.3. This theorem asserts that if the number of snapshots (i.e., $u^{p}(\bm{x},\bm{\xi}_{i}),\ i=1,\ldots,n$) is sufficiently large, the non-zero singular values of ${\mathbb{U}}^{{\mathrm{samp}}}_{D\times n}$ can be arbitrarily close to the singular values of $\mathbb{U}^{p}_{D\times n_{\xi}}$ with a scale factor. In other words, the range of $\mathbb{U}^{{\mathrm{samp}}}_{D\times n}$ aligns with the range of $\mathbb{U}^{p}_{D\times n_{\xi}}$. This suggests that the left singular vectors of $\mathbb{U}^{{\mathrm{samp}}}_{D\times n}$ can be used in place of the left singular vectors of $\mathbb{U}^{p}_{D\times n_{\xi}}$ as the basis functions of the physical space.
To further enhance efficiency, we employ a greedy algorithm with a residual-free indicator [31] to identify the most important $k$ sampling points from a set of random samples $\bm{\Theta}$. For computational efficiency, both snapshots and residual-free indicators are computed on a coarse grid. The identified $k$ sampling points are denoted as $\{\bm{\xi}_{1}^{{\mathrm{c}}},\ldots\bm{\xi}_{k}^{{\mathrm{c}}}\}$, and we then compute $u(\bm{x},\bm{\xi}_{i}^{{\mathrm{c}}})$ for $i=1,\ldots,k$ on the fine grid. These fine grid snapshots are denoted as $u^{{\mathrm{f}}}(\bm{x},\bm{\xi}_{i}^{{\mathrm{c}}})$. Suppose that the SVD decomposition of the quasimatrix $\mathbb{U}^{{\mathrm{f}}}$ is given by
$\mathbb{U}^{{\mathrm{f}}}=\widetilde{\mathbb{U}}\widetilde{\bm{\Sigma}}\widetilde{\bm{V}}^{*},$
where
$\mathbb{U}^{{\mathrm{f}}}=\left[\begin{array}[]{llll}\Big{|}&\Big{|}&&\Big{|}\\ u^{{\mathrm{f}}}(\bm{x},\bm{\xi}_{1}^{{\mathrm{c}}})&u^{{\mathrm{f}}}(\bm{x},\bm{\xi}_{2}^{{\mathrm{c}}})&\ldots&u^{{\mathrm{f}}}(\bm{x},\bm{\xi}_{k}^{{\mathrm{c}}})\\ \Big{|}&\Big{|}&&\Big{|}\end{array}\right].$
We then use
$X^{(k)}=\{\widetilde{u}_{1},\widetilde{u}_{2},\ldots,\widetilde{u}_{k}\}$
as the basis functions of the physical space, where $k$ is the same number as in (18), and $\widetilde{u}_{1},\widetilde{u}_{2},\ldots,\widetilde{u}_{k}$ are the columns of $\widetilde{\mathbb{U}}$.
Suppose that $\bm{W}^{(k)}$ and $\bm{V}^{(k)}$ are matrices satisfying
$\widetilde{\mathbb{U}}=\mathbb{X}\bm{W}^{(k)},\ \widetilde{\mathbb{S}}^{*}=(\bm{V}^{(k)})^{*}(\mathbb{S}^{p}_{\Gamma\times n_{\xi}})^{*},$
where
$\mathbb{X}=\left[\begin{array}[]{llll}\Big{|}&\Big{|}&&\Big{|}\\ \psi_{1}(\bm{x})&\psi_{2}(\bm{x})&\ldots&\psi_{n_{x}}(\bm{x})\\ \Big{|}&\Big{|}&&\Big{|}\end{array}\right],\ \widetilde{\mathbb{S}}^{*}=\left[\begin{array}[]{ccc}\text{---}&\widehat{\phi}_{1}&\text{---}\\ \text{---}&\widehat{\phi}_{2}&\text{---}\\ &\vdots&\\ \text{---}&\widehat{\phi}_{k}&\text{---}\end{array}\right],$
and the matrix $\bm{V}^{(k)}$ consists of the first $k$ columns of $\widehat{\bm{V}}$ defined in (16). The simultaneous low-rank projected matrix equation is then given by
$\sum_{i=1}^{K}\bm{A}_{i}^{(k)}\bm{X}^{(k)}(\bm{G}_{i}^{(k)})^{\mathrm{T}}=\bm{f}^{(k)}(\bm{h}^{(k)})^{\mathrm{T}},$ (19)
where
$\bm{A}_{i}^{(k)}$ $=(\bm{W}^{(k)})^{*}\bm{A}_{i}\bm{W}^{(k)},$ $\bm{f}^{(k)}$ $=(\bm{W}^{(k)})^{*}\bm{f},$ (20)
$\bm{G}_{i}^{(k)}$ $=(\bm{V}^{(k)})^{\mathrm{T}}\bm{G}_{i}\left((\bm{V}^{(k)})^{*}\right)^{\mathrm{T}},$ $\bm{h}^{(k)}$ $=(\bm{V}^{(k)})^{\mathrm{T}}\bm{h}.$ (21)
The low-rank projection approximation of the solution can then be given by
$u^{p}_{k}(\bm{x},\bm{\xi})=\widetilde{\mathbb{U}}\bm{X}^{(k)}\widetilde{\mathbb{S}}^{*}=\mathbb{X}\bm{W}^{(k)}\bm{X}^{(k)}(\bm{V}^{(k)})^{*}(\mathbb{S}^{p}_{\Gamma\times n_{\xi}})^{*}.$ (22)
In practical implementations, all operations are conducted in matrix form. In this study, the meshgrid in the physical space is uniform, and the $L^{2}$ norm of a function is approximated by the $2$-norm of the corresponding vector multiplied by a factor of $1/n_{x}$. Under these conditions, we employ matrix SVD to replace quasimatrix SVD decomposition. Algorithm 1 provides the pseudocode for the SimLRP approach.
Algorithm 1 A simultaneous low-rank projection (SimLRP) approach
1:INPUT: $p$: the gPC oder; $\mathtt{RelTol}$: the tolerance in(18); $T^{{\mathrm{c}}}$: a coarse physical grid; $T$: a fine physical grid; $\bm{\Theta}$: a set of random samples.
2:Generate $\bm{A}_{i}^{{\mathrm{c}}},\bm{f}^{{\mathrm{c}}}$ for the coarse physical gird $T^{{\mathrm{c}}}$, and $\bm{A}_{i},\bm{f}$ for the fine physical gird $T$, generate $\bm{G}_{i}$ and $\bm{h}$.
3:Solve the matrix equation
$\sum_{i=1}^{K}\bm{A}_{i}^{{\mathrm{c}}}\bm{U}^{{\mathrm{c}}}\bm{G}_{i}^{\mathrm{T}}=\bm{f}^{{\mathrm{c}}}\bm{h}^{\mathrm{T}},$
4:Compute SVD decomposition of $\bm{U}^{{\mathrm{c}}}$, i.e., $\bm{U}^{{\mathrm{c}}}=\widehat{\bm{U}}\widehat{\bm{\Sigma}}\widehat{\bm{V}}^{*}$.
5:$\ k$longleftarrow the smallest number satisfying (18), where $\widehat{\sigma}_{1},\ldots,\widehat{\sigma}_{n_{\xi}}$ are the diagonal of $\widehat{\bm{\Sigma}}$.
6:$\bm{V}^{(k)}$longleftarrow the first $k$ columns of $\widehat{\bm{V}}$.
7:$\ \bm{\theta}$longleftarrow the most important $k$ samples selected from $\bm{\Theta}$.
8:$\bm{U}^{{\mathrm{f}}}$longleftarrow snapshots on the fine physical grid corresponding to $\bm{\theta}$.
9:$\bm{W}^{(k)}$longleftarrow the first $k$ left singular vectors of $\bm{U}^{{\mathrm{f}}}$.
10:$\bm{X}^{(k)}$$\longleftarrow$ solve (19).
11:Construct low-rank approximation by (22).
3.3 An alternating low-rank projection approach
While the SimLRP approach provides the capability to simultaneously reduce the dimensionality of both the physical and stochastic spaces, the tolerance indicator (18) may deteriorate for smaller values of $\mathtt{RelTol}$. Drawing inspiration from [32, 33], we propose an alternating low-rank projection (AltLRP) approach to mitigate the deterioration of tolerance indicators for smaller values.
The idea is quite straightforward. After applying SimLRP, we alternatively implement the low-rank projection in the stochastic and physical spaces to refine and update the low-rank approximation. Suppose we have obtained $\bm{W}^{(k)}$, $\bm{X}^{(k)}$, $\bm{V}^{(k)}$ from the SimLRP approach. We then solve the following systems of equations in sequence:
$\sum_{i=1}^{K}\bm{A}_{i}\bm{Y}^{(k)}(\bm{G}_{i}^{(k)})^{\mathrm{T}}=\bm{f}(\bm{h}^{(k)})^{\mathrm{T}},$ (23)
$\sum_{i=1}^{K}\bm{A}_{i}^{(k)}\bm{Z}^{(k)}(\bm{G}_{i})^{\mathrm{T}}=\bm{f}^{(k)}(\bm{h})^{\mathrm{T}}.$ (24)
Note that once $\bm{Y}^{(k)}$ or $\bm{Z}^{(k)}$ is solved, $\bm{W}^{(k)}$ or $\bm{V}^{(k)}$, and thus $\bm{A}^{(k)}_{i}$ or $\bm{G}^{(k)}_{i}$, should be updated immediately. To reduce the number of iterations in solving (23) and (24), the solution from the previous step can be used to design initial guesses.
Algorithm 2 Alternating low-rank projection (AltLRP) approach
1:Implement Algorithm 1, and keep the notations.
[…]
3:for $i=1,2,\ldots i_{\max}$ do
4:     Update $\bm{G}_{i}^{(k)}$ and $\bm{h}^{(k)}$ by (21).
5:     $\bm{Y}^{(k)}$longleftarrow solve (23) with initial guess $\bm{W}^{(k)}\bm{\Sigma}^{(k)}$.
[…]
7:     Update $\bm{A}_{i}^{(k)}$ and $\bm{f}^{(k)}$ by (20).
8:     $\bm{Z}^{(k)}$longleftarrow solve (24) with initial guess $\bm{\Sigma}^{(k)}(\bm{V}^{(k)})^{*}$.
[…]
10:end for
11:Update $\bm{A}_{i}^{(k)}$, $\bm{f}^{(k)}$, $\bm{G}_{i}^{(k)}$, $\bm{h}^{(k)}$ by (20)–(21).
12:$\bm{X}^{(k)}$$\longleftarrow$ solve (19) with initial guess $\bm{\Sigma}^{(k)}$.
13:construct low-rank approximation by (22).
Algorithm 3 Truncated SVD decomposition
1:function [$\bm{W}^{(k)}$,$\bm{\Sigma}^{(k)}$, $\bm{V}^{(k)}$] = TruncatedSVD($\bm{X}$,$\bm{W}^{(k)}$, $\bm{V}^{(k)}$)
2:     Compute the SVD decomposition of $\bm{X}$, i.e., $\bm{X}=\bm{W}\bm{\Sigma}\bm{V}^{*}$.
3:     $\bm{W}^{(k)}_{{\mathrm{temp}}}$longleftarrow first $k$ columns of $\bm{W}$.
4:     $\bm{V}^{(k)}_{{\mathrm{temp}}}$longleftarrow first $k$ rows of $\bm{V}$.
5:     $\bm{\Sigma}^{(k)}_{{\mathrm{temp}}}$longleftarrow first $k\times k$ leading principal submatrix of $\bm{\Sigma}$.
6:     $\bm{W}^{(k)}$$\longleftarrow$ $\bm{W}^{(k)}\bm{W}^{(k)}_{{\mathrm{temp}}}$, $\bm{V}^{(k)}$$\longleftarrow$ $\bm{V}^{(k)}\bm{V}^{(k)}_{{\mathrm{temp}}}$, $\bm{\Sigma}^{(k)}$$\longleftarrow$ $\bm{\Sigma}^{(k)}_{{\mathrm{temp}}}$.
7:end function
Algorithm 2 provides the pseudocode for the AltLRP approach. Since $\bm{W}^{(k)}$, $\bm{X}^{(k)}$, $\bm{V}^{(k)}$ obtained from the SimLRP approach are already quite accurate, typically only one iteration (i.e., $i_{\max}=1$) is required in the AltLRP approach to achieve the desired accuracy.
4 Numerical results
In this section, we consider two problems: a diffusion problem and a Helmholtz problem. For the diffusion problem, we employ the standard finite element method to generate the matrices $\bm{A}_{i}$, where $i\in\{0,1,\ldots,K\}$. For the Helmholtz problem, we utilize the codes associated with [34] for convenience. All the results presented in this study are obtained using MATLAB R2016a on a desktop computer with 2.90GHz Intel Core i7-10700 CPU.
To assess accuracy, we define the relative error as follows:
$\mathtt{RelErr}=\frac{\left\|\bm{W}^{(k)}\bm{X}^{(k)}\bm{V}^{(k)}-\bm{U}\right\|_{F}}{\left\|\bm{U}\right\|_{F}},$ (25)
where $\bm{W}^{(k)}$, $\bm{X}^{(k)}$, $\bm{V}^{(k)}$ are the matrices corresponding to the low-rank approximation with rank $k$, and $\bm{U}$ is the matrix in (9).
4.1 Test problem 1
In this problem, we investigate the diffusion equation with random inputs, given by
$-\nabla\cdot(a(\bm{x},\bm{\xi})\nabla u(\bm{x},\bm{\xi}))=1&\quad\mbox{in}\quad D\times\Gamma,\\ u(\bm{x},\bm{\xi})=0&\quad\mbox{on}\quad\partial D\times\Gamma,\\ \end{aligned}\right.$
where $D=[-1,1]^{2}$ is the spatial domain, and $\partial D$ represents the boundary of $D$. The diffusion coefficient $a(\bm{x},\bm{\xi})$ is modeled as a truncated Karhunen–Loève (KL) expansion [6, 35] of a random field with a mean function $a_{0}(\bm{x})=0.2$, a standard deviation $\sigma=0.1$, and the covariance function $\mathrm{Cov}\,(\bm{x},\bm{y})$ given by
$\mathrm{Cov}\,(\bm{x},\bm{y})=\sigma^{2}\exp\left(-\frac{|x_{1}-y_{1}|}{c}-\frac{|x_{2}-y_{2}|}{c}\right),$
where $\bm{x}=[x_{1},x_{2}]^{T}$, $\bm{y}=[y_{1},y_{2}]^{T}$ and $c=1$ is the correlation length. The KL expansion takes the form
$a(\bm{x},\bm{\xi})=a_{0}(\bm{x})+\sum_{i=1}^{N}a_{i}(\bm{x})\xi_{i}=a_{0}(\bm{x})+\sum_{i=1}^{N}\sqrt{\lambda_{i}}c_{i}(\bm{x})\xi_{i},$
where $\{\lambda_{i},c_{i}(\bm{x})\}_{i=1}^{N}$ are the eigenpairs of $\mathrm{Cov}\,(\bm{x},\bm{y})$, $\{\xi_{i}\}^{N}_{i=1}$ are uncorrelated random variables, and $N$ is the number of KL modes retained. For this test problem, we assume that the random variables $\{\xi_{i}\}^{N}_{i=1}$ are identically independent distributed uniform random variables on $[-1,1]$.
In this test problem, the gPC order $p$ is set to $4$, the degrees of freedom (DOF) of the coarse grid in physical space (i.e., $\bm{T}^{c}$) is set to $n_{x}=33\times 33$, and the cardinality of the random set $\bm{\Theta}$ is set to $|\bm{\Theta}|=\min(2n_{\xi},n_{x})$. All the linear systems, i.e., (19) and (23)-(24), are solved using the preconditioned conjugate gradients (PCG) method with a mean-based preconditioner [15] and a tolerance of $10^{-8}$.
(a) $N=5$
(b) $N=7$
(c) $N=10$
Figure 1: Singular values for different $N$.
In this figure, ‘fine’ refers to the singular values of $\mathbb{U}^{p}_{D\times n_{\xi}}$ with $n_{x}=257^{2}$; ‘coarse’ refers to the singular values of $\mathbb{U}^{{\mathrm{c}}}_{D\times n_{\xi}}$ with $n_{x}=33^{2}$; and ‘random’ refers to the singular values of $\widehat{\mathbb{U}}^{{\mathrm{samp}}}_{D\times n}$ in Corollary 3.3.1 with $|\bm{\Theta}|=\min(2n_{\xi},n_{x})$ and $n_{x}=33^{2}$.
In Figure 1, we plot the singular values for different $N$. The DOF of the physical space is $n_{x}=33^{2}$ for the coarse grid and $n_{x}=257^{2}$ for the fine grid. It is noteworthy that, in this example, the $L^{2}$ norm of a function is simply approximated by the $2$-norm of its corresponding vector, multiplied by a factor of $1/n_{x}$, since the mesh grid in physical space is uniform. From the figure, we observe that the singular values for ‘fine’, ‘coarse’ and ‘random’ coincide very well with each other, reflecting the conclusion of Theorem 3.3. It also demonstrates that the singular values of $\mathbb{U}^{p}_{D\times n_{\xi}}$ are not very sensitive to the DOF of the physical mesh grid, and thus allowing the use of singular values computed on a coarse grid to determine the rank $k$ for a desired accuracy $\mathtt{RelTol}$.
(a) $N=5,\ n_{x}=129^{2}$
(b) $N=7,\ n_{x}=129^{2}$
(c) $N=10,\ n_{x}=129^{2}$
(d) $N=5,\ n_{x}=257^{2}$
(e) $N=7,\ n_{x}=257^{2}$
(f) $N=10,\ n_{x}=257^{2}$
Figure 2: Relative errors with respect to different $\mathtt{RelTol}$.
In Figure 2, we present the relative errors concerning different values of $\mathtt{RelTol}$, where the relative error is defined by (25). The ‘optimal’ solution is obtained through the SVD decomposition of $\bm{U}$ with the corresponding $k$ terms retained. In the AltLRP approach, the value of $i_{\max}$ is set to be $1$. From the figure, it is evident that the AltLRP approach performs nearly optimally. While the SimLRP approach also demonstrates good performance, the tolerance indicator (18) for it deteriorates for smaller values of $\mathtt{RelTol}$.
(a) $N=5,\ n_{x}=129^{2}$
(b) $N=7,\ n_{x}=129^{2}$
(c) $N=10,\ n_{x}=129^{2}$
(d) $N=5,\ n_{x}=257^{2}$
(e) $N=7,\ n_{x}=257^{2}$
(f) $N=10,\ n_{x}=257^{2}$
Figure 3: Relative errors with respect to CPU time in seconds.
In Figure 3, we present the relative errors with respect to CPU time in seconds. The figure reveals that, for a comparable level of accuracy, the CPU time of the SimLRP approach is slightly smaller than that of the AltLRP approach when $N=5$ or when the relative error is large. However, for $N=7$ and $N=10$, AltLRP outperforms SimLRP. This observation suggests that, for low-accuracy approximations or low-dimensional problems, the SimLRP approach may be preferred, whereas for higher accuracy or higher dimensional problems, the AltLRP approach is the better choice.
Table 1: CPU time in seconds for standard stochastic Galerkin method.
$n_{x}$ $N=5$ $N=7$ $N=10$
$129^{2}$ $38.38$ $109.37$ $394.43$
$257^{2}$ $229.70$ $659.54$ $2352.73$
Table 1 displays the CPU time in seconds for the standard stochastic Galerkin method. From Figure 3 and Table 1, it is evident that both the proposed methods are more efficient than the standard stochastic Galerkin method for achieving certain levels of accuracy.
4.2 Test problem 2
In this test problem, we consider the stochastic Helmholtz problem given by
$\nabla^{2}u+a^{2}(\bm{x},\bm{\xi})u=f(\bm{x})\quad\mbox{in}\quad D\times\Gamma,$
with the Sommerfeld radiation boundary condition. Here, $D=[0,1]^{2}$ is the domain of interest, and the Helmholtz coefficient $a(\bm{x},\bm{\xi})$ is a truncated KL expansion of a random field with a mean function $a_{0}(\bm{x})=4\cdot(2\pi)$, a standard deviation $\sigma=0.8\pi$, and the covariance function $\mathrm{Cov}\,(\bm{x},\bm{y})$ given by
$\mathrm{Cov}\,(\bm{x},\bm{y})=\sigma^{2}\exp\left(-\frac{|x_{1}-y_{1}|}{c}-\frac{|x_{2}-y_{2}|}{c}\right),$
where $\bm{x}=[x_{1},x_{2}]^{T}$, $\bm{y}=[y_{1},y_{2}]^{T}$ and $c=4$ is the correlation length. Note that the KL expansion takes the form
$a(\bm{x},\bm{\xi})=a_{0}(\bm{x})+\sum_{i=1}^{N}a_{i}(\bm{x})\xi_{i}=a_{0}(\bm{x})+\sum_{i=1}^{N}\sqrt{\lambda_{i}}c_{i}(\bm{x})\xi_{i},$
where $\{\lambda_{i},c_{i}(\bm{x})\}_{i=1}^{N}$ are the eigenpairs of $\mathrm{Cov}\,(\bm{x},\bm{y})$, $\{\xi_{i}\}^{N}_{i=1}$ are uncorrelated random variables, and $N$ is the number of KL modes retained. The Gaussian point source at the center of the domain is used as the source term, i.e.,
$f(\bm{x})=\mbox{e}^{-(8\cdot 4)^{2}((x_{1}-0.5)^{2}+(x_{2}-0.5)^{2})}.$
For this test problem, we assume that the random variables $\{\xi_{i}\}^{N}_{i=1}$ are independent and uniformly distributed within the range $[-1,1]$. We use the perfectly matched layers (PML) to simulate the Sommerfeld condition [36], and generate the matrices $\{\bm{A}_{i}\}_{i=1}^{K}$ using the codes associated with [34].
In this test problem, the gPC order $p$ is set to $4$, the degrees of freedom (DOF) of the coarse grid in physical space (i.e., $\bm{T}^{c}$) is set to $n_{x}=33\times 33$, and the cardinality of the random set $\bm{\Theta}$ is set to $|\bm{\Theta}|=\min(2n_{\xi},n_{x})$. All the linear systems, i.e., (19) and (23)-(24), are solved using the preconditioned bi-conjugate gradient stabilized (Bi-CGSTAB) method with a mean-based preconditioner [15] and a tolerance of $10^{-8}$.
(a) $N=5$
(b) $N=7$
(c) $N=10$
Figure 4: Singular values for different $m$ and $\mathtt{RelTol}$.
In this figure, ‘fine’ refers to the singular values of $\mathbb{U}^{p}_{D\times n_{\xi}}$ with $n_{x}=257^{2}$; ‘coarse’ refers to the singular values of $\mathbb{U}^{{\mathrm{c}}}_{D\times n_{\xi}}$ with $n_{x}=33^{2}$; and ‘random’ refers to the singular values of $\widehat{\mathbb{U}}^{{\mathrm{samp}}}_{D\times n}$ in Corollary 3.3.1 with $|\bm{\Theta}|=\min(2n_{\xi},n_{x})$ and $n_{x}=33^{2}$.
In Figure 4, we plot the singular values for different $N$. The DOF of the physical space is $n_{x}=33^{2}$ for the coarse grid and $n_{x}=257^{2}$ for the fine grid. It is noteworthy that, in this example, the $L^{2}$ norm of a function is simply approximated by the $2$-norm of its corresponding vector, multiplied by a factor of $1/n_{x}$, since the mesh grid in physical space is uniform. From the figure, we observe that the singular values for ‘fine’, ‘coarse’ and ‘random’ coincide very well with each other, reflecting the conclusion of Theorem 3.3. It also demonstrates that the singular values of $\mathbb{U}^{p}_{D\times n_{\xi}}$ are not very sensitive to the DOF of the physical mesh grid, and thus allowing the use of singular values computed on a coarse grid to determine the rank $k$ for a desired accuracy $\mathtt{RelTol}$.
(a) $N=5,\ n_{x}=129^{2}$
(b) $N=7,\ n_{x}=129^{2}$
(c) $N=10,\ n_{x}=129^{2}$
(d) $N=5,\ n_{x}=257^{2}$
(e) $N=7,\ n_{x}=257^{2}$
(f) $N=10,\ n_{x}=257^{2}$
Figure 5: Relative errors with respect to different $\mathtt{RelTol}$.
In Figure 5, we present the relative errors concerning different values of $\mathtt{RelTol}$, where the relative error is defined by (25). The ‘optimal’ solution is obtained through the SVD decomposition of $\bm{U}$ with the corresponding $k$ terms retained. In the AltLRP approach, the value of $i_{\max}$ is set to be $1$. From the figure, it is evident that the AltLRP approach performs nearly optimally. While the SimLRP approach also demonstrates good performance, the tolerance indicator (18) for it deteriorates for smaller values of $\mathtt{RelTol}$.
(a) $N=5,\ n_{x}=129^{2}$
(b) $N=7,\ n_{x}=129^{2}$
(c) $N=10,\ n_{x}=129^{2}$
(d) $N=5,\ n_{x}=257^{2}$
(e) $N=7,\ n_{x}=257^{2}$
(f) $N=10,\ n_{x}=257^{2}$
Figure 6: Relative errors with respect to CPU time in seconds.
In Figure 6, we present the relative errors with respect to CPU time in seconds. The figure reveals that, for a comparable level of accuracy, the CPU time of the SimLRP approach is slightly smaller than that of the AltLRP approach when $N=5$ or when the relative error is large. However, for $N=7$ and $N=10$, AltLRP outperforms SimLRP. This observation suggests that, for low-accuracy approximations or low-dimensional problems, the SimLRP approach may be preferred, whereas for higher accuracy or higher dimensional problems, the AltLRP approach is the better choice.
Table 2: CPU time in seconds for standard stochastic Galerkin method.
$n_{x}$ $N=5$ $N=7$ $N=10$
$129^{2}$ $57.74$ $214.71$ $908.00$
$257^{2}$ $308.87$ $1138.43$ $4645.46$
Table 2 displays the CPU time in seconds for the standard stochastic Galerkin method. Again, from Figure 6 and Table 2, it is evident that both the proposed methods are more efficient than the standard stochastic Galerkin method for achieving certain levels of accuracy.
5 Conclusion
In this work, we develop both a simultaneous low-rank projection (SimLRP) approach and an alternating low-rank projection (AltLRP) approach to compute the low-rank approximation of the solution for PDEs with random inputs. To identify the rank required for a desired accuracy, we propose a systematic strategy grounded in theoretical analysis, exploring the singular values of quasimatrices associated with truncated generalized polynomial chaos (gPC) expansions of the solution. Both the proposed methods exhibit enhanced computational efficiency compared to the conventional stochastic Galerkin method, particularly when dealing with numerous physical degrees of freedom. Numerical results reveal that SimLRP excels in addressing low-dimensional problems, whereas AltLRP demonstrates superior performance for moderate-dimensional problems. In the future, our focus will be on developing effective algorithms tailored for high-dimensional problems utilizing both AltLRP and SimLRP.